OpenStudy (anonymous):

Jocelyn writes the equation of a line that passes through the point (4,0) and that is perpendicular to the line y= -2X + 1. IF Jocelyn write her equation in the form y= MX +B, what is the value of B?

jimthompson5910 (jim_thompson5910):

the slope of y= -2x + 1 is -2

jimthompson5910 (jim_thompson5910):

flip the fraction and flip the sign to get m = 1/2 this is the perpendicular slope

jimthompson5910 (jim_thompson5910):

y = mx + b y = (1/2)x + b ... plug in the perpendicular slope 0 = (1/2)*(4) + b ... plug in the given point (4,0) 0 = 4/2 + b keep going to solve for b

jimthompson5910 (jim_thompson5910):

0 = 4/2 + b turns into 0 = 2 + b ... ... b = ??